a number is subtracted from 20 and from 17.the product of the numbers so obtained is 180.find the original number
Answer by ikleyn(52864)

.
(20-x)*(17-x) = 180.

180 = 12*15  is an appropriate factorization of  180  into the product of two numbers that differs in  3.

So,  20-x = 15  (greater of these two factors),

which implies

x = 5.

180 = (-12)*(-15)  is also an appropriate factorization of  180  into the product of two numbers that differs in  3.

So,  20-x = -12,

which implies

x = 32.

Answer.  Two solutions are  x=5  and  x=32.

P.S.  Surely,  you can solve the quadratic equation above by using the quadratic formula.